# A unified relative entropy framework for macroscopic limits of Vlasov–Fokker–Planck equations

### v1

@@ 4.2. High-field limit
which guarantees the desired convergence of $\rho^{\varepsilon}$ and $f^{\varepsilon}$ in BL-distance. This completes the proof.
@@ Appendix A Formal derivation of the gSQG limit
$\varepsilon\partial_{t}f^{\varepsilon}+\xi\cdot\nabla f^{\varepsilon}=\nabla_{\xi}\cdot(\nabla_{\xi}f^{\varepsilon}+\xi f^{\varepsilon})+\frac{1}{\varepsilon}\,\xi^{\perp}\cdot\nabla_{\xi}f^{\varepsilon}-\nabla(-\Delta)^{-\alpha}\rho_{f^{\varepsilon}}\cdot\nabla_{\xi}f^{\varepsilon},$
@@
We consider a Hilbert expansion $f^{\varepsilon}=f^{0}+\varepsilon f^{1}+O(\varepsilon^{2})$. The leading constraint $L_{-1}f^{0}=0$ enforces radial symmetry in $\xi$. Accordingly, we set
@@
$L_{-1}f^{1}+\xi\cdot\nabla f^{0}-\nabla\Phi^{0}\cdot\nabla_{\xi}f^{0}+\nabla_{\xi}\cdot(\nabla_{\xi}f^{0}+\xi f^{0})=0,\quad\Phi^{0}:=(-\Delta)^{-\alpha}\rho^{0},$
@@
$L_{-1}f^{1}+(\xi\cdot\nabla\rho^{0})M_{1,0}+(\rho^{0}\nabla\Phi^{0}\cdot\xi)M_{1,0}=0.$
@@
$\Pi(\xi\cdot\nabla f^{0}-\nabla\Phi\cdot\nabla_{\xi}f^{0})=0.$
On the complementary subspace $(I-\Pi)L^{2}_{\xi}$, $L_{-1}$ is invertible, so that the non-radial part of the right-hand side uniquely determines $f^{1}$ up to a radial component. Using $L_{-1}\big((\xi\cdot a)M_{1,0}\big)=(\xi^{\perp}\cdot a)M_{1,0}$ for any constant vector $a$, we may choose the corrector
$f^{1}=-\left((\nabla\rho^{0}+\rho^{0}\nabla\Phi^{0})^{\perp}\cdot\xi\right)\,M_{1,0},$
@@
$m^{1}(t,x):=\int_{\mathbb{R}^{2}}\xi f^{1}\,d\xi=-\left(\nabla\rho^{0}+\rho^{0}\nabla\Phi^{0}\right)^{\perp}\int_{\mathbb{R}^{2}}\xi\otimes\xi\,M_{1,0}\,d\xi=-\nabla^{\perp}\rho^{0}-\rho^{0}\nabla^{\perp}\Phi^{0}.$
Inserting this into the continuity equation $\partial_{t}\rho^{0}+\nabla\cdot m^{1}=0$ yields
$\partial_{t}\rho^{0}-\nabla\cdot(\rho^{0}\nabla^{\perp}\Phi^{0})+\nabla\cdot(\nabla^{\perp}\rho^{0})=0,$
and here the divergence-free identity $\nabla\cdot\nabla^{\perp}\equiv 0$ in two dimensions produces an exact cancellation of the intermediate $\nabla^{\perp}\rho^{0}$ term, so
@@
which coincides with (1.12). We stress that a term involving $\nabla^{\perp}\rho^{0}$ genuinely appears in the kinetic corrector and in the intermediate flux. It disappears only after applying the divergence in the continuity equation, thanks to the divergence-free identity $\nabla\cdot\nabla^{\perp}=0$ in two dimensions. This cancellation explains why moment equations alone are not sufficient to capture the correct structure, and why the fast rotation averaging coupled with the projection $\Pi$ provides the natural framework for the gSQG limit.
Appendix B Propagation of regularity for the limit equations

### v2

@@ 4.2. High-field limit
which gives the desired quantitative estimate for $\rho^{\varepsilon}$. This completes the proof.
Remark 4.2.
Recall ${\rm u}_{\varepsilon}:=-\varepsilon\left(\nabla^{\perp}\log\rho+\nabla^{\perp}(-\Delta)^{-\alpha}\rho\right)$. Then Corollary 3.1, together with (4.15) and the above $H^{-\gamma}$ estimate for $\rho^{\varepsilon}-\rho$, yields
$\int_{0}^{T}\|f-M_{\rho,{\rm u}_{\varepsilon}}\|_{H_{x,\xi}^{-\sigma}}^{2}\,dt$ $\leq C\int_{0}^{T}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|\nabla_{\xi}f^{\varepsilon}+(\xi-u^{\varepsilon})f^{\varepsilon}|^{2}\,dxd\xi dt+C\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u}_{\varepsilon}|^{2}\,dxdt$
$\quad+C\|{\rm u}_{\varepsilon}\|_{L^{\infty}((0,T)\times\mathbb{R}^{d})}^{2}\int_{0}^{T}\|\rho^{\varepsilon}-\rho\|_{L^{1}}^{2}\,dt+C\int_{0}^{T}\|\rho^{\varepsilon}-\rho\|_{H^{-\sigma}}^{2}\,dt$
$\leq C\left(\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\gamma}}+\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon\right)^{2}$
for any $\sigma\geq\gamma$. In particular, the shifted Maxwellian $M_{\rho,{\rm u}_{\varepsilon}}$ provides a rigorously justified refined approximation of the kinetic distribution in a weak topology.
@@ Appendix A Formal derivation of the gSQG limit
$\varepsilon\partial_{t}f^{\varepsilon}+\xi\cdot\nabla f^{\varepsilon}=\nabla_{\xi}\cdot(\nabla_{\xi}f^{\varepsilon}+\xi f^{\varepsilon})+\nabla(-\Delta)^{-\alpha}\rho^{\varepsilon}\cdot\nabla_{\xi}f^{\varepsilon}-\frac{1}{\varepsilon}\,\xi^{\perp}\cdot\nabla_{\xi}f^{\varepsilon},$
@@
We consider a Hilbert expansion $f^{\varepsilon}=f^{0}+\varepsilon f^{1}+O(\varepsilon^{2})$. The leading constraint $L_{-1}f^{0}=0$ enforces radial symmetry in $\xi$. Accordingly, we set an ansatz for $f_{0}$ as
@@
$L_{-1}f^{1}+\xi\cdot\nabla f^{0}-\nabla(-\Delta)^{-\alpha}\rho^{0}\cdot\nabla_{\xi}f^{0}-\nabla_{\xi}\cdot(\nabla_{\xi}f^{0}+\xi f^{0})=0,$
@@
$L_{-1}f^{1}=(\xi\cdot\nabla\rho^{0})M_{1,0}+(\rho^{0}\nabla(-\Delta)^{-\alpha}\rho^{0}\cdot\xi)M_{1,0}.$
Equivalently,
$L_{-1}f^{1}=\bigl(\xi\cdot(\nabla\rho^{0}+\rho^{0}\nabla(-\Delta)^{-\alpha}\rho^{0})\bigr)M_{1,0}.$
@@
$\Pi\bigl(\xi\cdot\nabla f^{0}-\nabla(-\Delta)^{-\alpha}\rho^{0}\cdot\nabla_{\xi}f^{0}\bigr)=0,$
which actually holds for $f_{0}$ we set above. On the complementary subspace $(I-\Pi)L^{2}_{\xi}$, $L_{-1}$ is invertible, so that the non-radial part of the right-hand side uniquely determines $f^{1}$ up to a radial component. Using
$L_{-1}\big((\xi\cdot a)M_{1,0}\big)=(\xi^{\perp}\cdot a)M_{1,0}$
for any constant vector $a$, we may choose the corrector
$f^{1}=-\bigl((\nabla\rho^{0}+\rho^{0}\nabla(-\Delta)^{-\alpha}\rho^{0})^{\perp}\cdot\xi\bigr)\,M_{1,0},$
@@
$m^{1}(t,x):=\int_{\mathbb{R}^{2}}\xi f^{1}\,d\xi=-(\nabla\rho^{0}+\rho^{0}\nabla(-\Delta)^{-\alpha}\rho^{0})^{\perp}\int_{\mathbb{R}^{2}}\xi\otimes\xi\,M_{1,0}\,d\xi=-\nabla^{\perp}\rho^{0}-\rho^{0}\nabla^{\perp}(-\Delta)^{-\alpha}\rho^{0}.$
Inserting this into the continuity equation
$\partial_{t}\rho^{0}+\nabla\cdot m^{1}=0$
yields
$\partial_{t}\rho^{0}-\nabla\cdot(\nabla^{\perp}\rho^{0})-\nabla\cdot(\rho^{0}\nabla^{\perp}(-\Delta)^{-\alpha}\rho^{0})=0.$
Since $\nabla\cdot\nabla^{\perp}\equiv 0$ in two dimensions, the intermediate term involving $\nabla^{\perp}\rho^{0}$ vanishes, and therefore
@@
We stress that a term involving $\nabla^{\perp}\rho^{0}$ genuinely appears in the kinetic corrector and in the intermediate flux. It disappears only after applying the divergence in the continuity equation, thanks to the identity $\nabla\cdot\nabla^{\perp}=0$ in two dimensions. This cancellation explains why moment equations alone are not sufficient to capture the correct structure, and why the fast rotation averaging coupled with the projection $\Pi$ provides the natural framework for the gSQG limit.
Appendix B Formal optimal convergence rates
In this appendix, we discuss the formal convergence rates suggested by the asymptotic expansions in the three singular regimes considered in the paper.
B.1. Diffusive limit
We first consider the diffusive scaling
$\varepsilon\partial_{t}f^{\varepsilon}+\xi\cdot\nabla f^{\varepsilon}-\nabla(-\Delta)^{-\alpha}\rho^{\varepsilon}\cdot\nabla_{\xi}f^{\varepsilon}=\frac{1}{\varepsilon}\nabla_{\xi}\cdot(\nabla_{\xi}f^{\varepsilon}+\xi f^{\varepsilon}),$
where $\rho^{\varepsilon}=\int_{\mathbb{R}^{d}}f^{\varepsilon}\,d\xi$. We seek a formal expansion of the form
$f^{\varepsilon}=f_{0}+\varepsilon f_{1}+\varepsilon^{2}f_{2}+\cdots.$
Substituting this ansatz into the equation and comparing powers of $\varepsilon$, the leading-order term gives
$\nabla_{\xi}\cdot(\nabla_{\xi}f_{0}+\xi f_{0})=0.$
Hence $f_{0}$ lies in the kernel of the Fokker–Planck operator, and thus must be of the form
$f_{0}=M_{\rho,0}=\frac{\rho(t,x)}{(2\pi)^{\frac{d}{2}}}e^{-\frac{|\xi|^{2}}{2}}.$
To determine the first-order correction, we collect the terms of order $O(1)$:
$\xi\cdot\nabla f_{0}-\nabla(-\Delta)^{-\alpha}\rho\cdot\nabla_{\xi}f_{0}=\nabla_{\xi}\cdot(\nabla_{\xi}f_{1}+\xi f_{1}).$
Using
$\nabla f_{0}=M_{1,0}\nabla\rho,\quad\nabla_{\xi}f_{0}=-\xi M_{\rho,0},$
we obtain
$\xi\cdot\nabla f_{0}-\nabla(-\Delta)^{-\alpha}\rho\cdot\nabla_{\xi}f_{0}=\xi\cdot\nabla\rho\,M_{1,0}+\rho\,\xi\cdot\nabla(-\Delta)^{-\alpha}\rho\,M_{1,0}.$
Equivalently,
$\xi\cdot\nabla f_{0}-\nabla(-\Delta)^{-\alpha}\rho\cdot\nabla_{\xi}f_{0}=\xi\cdot\left(\nabla\rho+\rho\nabla(-\Delta)^{-\alpha}\rho\right)M_{1,0}.$
Since
$\nabla_{\xi}\cdot\left(\nabla_{\xi}(\xi_{j}M_{1,0})+\xi(\xi_{j}M_{1,0})\right)=-\xi_{j}M_{1,0},$
it follows formally that
$f_{1}=-\xi\cdot\left(\nabla\rho+\rho\nabla(-\Delta)^{-\alpha}\rho\right)M_{1,0}.$
Recalling that
$\nabla\rho+\rho\nabla(-\Delta)^{-\alpha}\rho=\rho\left(\nabla\log\rho+\nabla(-\Delta)^{-\alpha}\rho\right),$
we may also write
$f_{1}=-\rho\,\xi\cdot\left(\nabla\log\rho+\nabla(-\Delta)^{-\alpha}\rho\right)M_{1,0}.$
A crucial observation is that
$\int_{\mathbb{R}^{d}}f_{1}\,d\xi=0,$
since $f_{1}$ is odd in $\xi$. Therefore, the first-order correction does not contribute to the density variable. As a consequence, the expansion suggests the improved density approximation
$\rho^{\varepsilon}=\rho+O(\varepsilon^{2}).$
On the other hand, the momentum satisfies
$m^{\varepsilon}:=\int_{\mathbb{R}^{d}}\xi f^{\varepsilon}\,d\xi=\varepsilon\int_{\mathbb{R}^{d}}\xi f_{1}\,d\xi+O(\varepsilon^{2}),$
and hence
$\frac{1}{\varepsilon}m^{\varepsilon}=-\rho\left(\nabla\log\rho+\nabla(-\Delta)^{-\alpha}\rho\right)+O(\varepsilon).$
We are thus led to the formal convergence rates
$f^{\varepsilon}-M_{\rho,0}=O(\varepsilon),\quad\rho^{\varepsilon}-\rho=O(\varepsilon^{2}),$
and
$\frac{1}{\varepsilon}m^{\varepsilon}+\rho\Bigl(\nabla(-\Delta)^{-\alpha}\rho+\nabla\log\rho\Bigr)=O(\varepsilon).$
In particular, the density variable gains one order compared to the kinetic distribution itself. This additional order is due to the oddness of the first-order corrector in the velocity variable and is consistent with the parabolic structure of the diffusive limit. The estimates obtained in Theorem 1.1, especially the quantitative control in a negative Sobolev topology, are therefore compatible with the formally optimal diffusive scaling.
Concerning the optimal rate toward the local Maxwellian state $M_{\rho,{\rm u}_{\varepsilon}}$, recall that
${\rm u}_{\varepsilon}=-\varepsilon\left(\nabla(-\Delta)^{-\alpha}\rho+\nabla\log\rho\right).$
By the expression of $f_{1}$, we have
$\varepsilon f_{1}=\xi\cdot{\rm u}_{\varepsilon}\,M_{\rho,0}.$
Hence,
$f^{\varepsilon}=M_{\rho,0}+\xi\cdot{\rm u}_{\varepsilon}\,M_{\rho,0}+O(\varepsilon^{2}).$
On the other hand, since
$M_{\rho,{\rm u}}(\xi)=\rho M_{1,{\rm u}}(\xi),$
we apply the Taylor expansion with integral remainder with respect to $\rm u$} around ${\rm u}=0$. We first write
$M_{\rho,{\rm u}_{\varepsilon}}-M_{\rho,0}=\int_{0}^{1}\frac{d}{d\tau}M_{\rho,\tau{\rm u}_{\varepsilon}}\,d\tau=\rho\int_{0}^{1}M_{1,\tau{\rm u}_{\varepsilon}}(\xi-\tau{\rm u}_{\varepsilon})\cdot{\rm u}_{\varepsilon}\,d\tau.$
Subtracting the first-order term $\xi\cdot{\rm u}_{\varepsilon}\,M_{\rho,0}$, we obtain
$M_{\rho,{\rm u}_{\varepsilon}}-M_{\rho,0}-\xi\cdot{\rm u}_{\varepsilon}\,M_{\rho,0}=\rho(\xi\cdot{\rm u}_{\varepsilon})\int_{0}^{1}\bigl(M_{1,\tau{\rm u}_{\varepsilon}}-M_{1,0}\bigr)\,d\tau-\rho|{\rm u}_{\varepsilon}|^{2}\int_{0}^{1}\tau M_{1,\tau{\rm u}_{\varepsilon}}\,d\tau.$
Moreover, applying the fundamental theorem of calculus to the map $\nu\mapsto M_{1,\nu\tau{\rm u}_{\varepsilon}}$, we get
$M_{1,\tau{\rm u}_{\varepsilon}}-M_{1,0}=\int_{0}^{1}M_{1,\nu\tau{\rm u}_{\varepsilon}}(\xi-\nu\tau{\rm u}_{\varepsilon})\cdot\tau{\rm u}_{\varepsilon}\,d\nu.$
This gives
$|M_{\rho,{\rm u}_{\varepsilon}}-M_{\rho,0}-\xi\cdot{\rm u}_{\varepsilon}\,M_{\rho,0}|$ $\leq C\rho|\xi|\,|{\rm u}_{\varepsilon}|\int_{0}^{1}\int_{0}^{1}M_{1,\nu\tau{\rm u}_{\varepsilon}}\bigl(|\xi|+\nu\tau|{\rm u}_{\varepsilon}|\bigr)\tau|{\rm u}_{\varepsilon}|\,d\nu d\tau+C\rho|{\rm u}_{\varepsilon}|^{2}\int_{0}^{1}M_{1,\tau{\rm u}_{\varepsilon}}\,d\tau$
$\leq C|{\rm u}_{\varepsilon}|^{2}(1+|\xi|^{2})M_{\rho,0},$
at least formally for $|{\rm u}_{\varepsilon}|$ small, where we used that the shifted Maxwellians $M_{1,\nu\tau{\rm u}_{\varepsilon}}$ remain uniformly comparable to $M_{1,0}$. Hence,
$M_{\rho,{\rm u}_{\varepsilon}}=M_{\rho,0}+\xi\cdot{\rm u}_{\varepsilon}\,M_{\rho,0}+O\bigl(|{\rm u}_{\varepsilon}|^{2}(1+|\xi|^{2})M_{\rho,0}\bigr).$
Since ${\rm u}_{\varepsilon}=O(\varepsilon)$, it follows that
$M_{\rho,{\rm u}_{\varepsilon}}=M_{\rho,0}+\xi\cdot{\rm u}_{\varepsilon}\,M_{\rho,0}+O\bigl(\varepsilon^{2}(1+|\xi|^{2})M_{\rho,0}\bigr),$
and subsequently,
$f^{\varepsilon}-M_{\rho,{\rm u}_{\varepsilon}}=O\bigl(\varepsilon^{2}(1+|\xi|^{2})M_{\rho,0}\bigr).$
B.2. High-field limit
We next consider the high-field scaling
$\varepsilon\partial_{t}f^{\varepsilon}+\varepsilon\xi\cdot\nabla f^{\varepsilon}-\nabla(-\Delta)^{-\alpha}\rho^{\varepsilon}\cdot\nabla_{\xi}f^{\varepsilon}=\nabla_{\xi}\cdot(\nabla_{\xi}f^{\varepsilon}+\xi f^{\varepsilon}).$
We again assume an expansion
$f^{\varepsilon}=f_{0}+\varepsilon f_{1}+\varepsilon^{2}f_{2}+\cdots,\quad\rho^{\varepsilon}=\rho+\varepsilon\rho_{1}+\varepsilon^{2}\rho_{2}+\cdots.$
At leading order, the equation becomes
$\nabla(-\Delta)^{-\alpha}\rho\cdot\nabla_{\xi}f_{0}=\nabla_{\xi}\cdot(\nabla_{\xi}f_{0}+\xi f_{0}).$
This relation characterizes the local equilibrium associated with the drift velocity
$u=-\nabla(-\Delta)^{-\alpha}\rho,$
namely
$f_{0}=M_{\rho,u}=M_{\rho,-\nabla(-\Delta)^{-\alpha}\rho}.$
Indeed, one may rewrite the leading-order equation as
$\nabla_{\xi}\cdot\left(\nabla_{\xi}f_{0}+(\xi-u)f_{0}\right)=0,$
whose equilibrium is exactly the Maxwellian centered at $u$. Thus, the dominant balance is no longer around the zero-velocity Maxwellian, but around the shifted local Maxwellian determined by the interaction field.
At the next order, we obtain
$\partial_{t}f_{0}+\xi\cdot\nabla f_{0}-\nabla(-\Delta)^{-\alpha}\rho\cdot\nabla_{\xi}f_{1}-\nabla(-\Delta)^{-\alpha}\rho_{1}\cdot\nabla_{\xi}f_{0}=\nabla_{\xi}\cdot(\nabla_{\xi}f_{1}+\xi f_{1}).$
Equivalently, introducing the linearized Fokker–Planck operator around the shifted equilibrium
$\mathcal{L}_{u}g:=\nabla_{\xi}\cdot\left(\nabla_{\xi}g+(\xi-u)g\right),$
the first-order corrector $f_{1}$ is formally determined by
$\mathcal{L}_{u}f_{1}=\partial_{t}f_{0}+\xi\cdot\nabla f_{0}-\nabla(-\Delta)^{-\alpha}\rho_{1}\cdot\nabla_{\xi}f_{0}.$
Hence, unlike in the diffusive regime, the first-order correction can still be identified at the formal level, but its structure is more involved because the reference equilibrium already carries the nontrivial drift $u=-\nabla(-\Delta)^{-\alpha}\rho$.
In particular, there is no reason to expect that
$\int_{\mathbb{R}^{d}}f_{1}\,d\xi=\rho_{1}$
vanishes in general. Therefore, unlike in the diffusive regime, the density variable does not benefit from an additional cancellation at first order. For this reason, the formal expansion suggests only
$f^{\varepsilon}=M_{\rho,-\nabla(-\Delta)^{-\alpha}\rho}+O(\varepsilon),$
and, at the macroscopic level,
$\rho^{\varepsilon}-\rho=O(\varepsilon),\quad m^{\varepsilon}+\rho\nabla(-\Delta)^{-\alpha}\rho=O(\varepsilon).$
Accordingly, one expects first-order convergence for both the kinetic distribution and its macroscopic moments. This is consistent with the quantitative estimates established in Theorem 1.2.
B.3. gSQG limit
We finally consider the strong magnetic field regime leading to the gSQG equation. Since the formal derivation of the limiting equation has already been presented in Appendix A, we only recall here the main consequence relevant to the present appendix, namely the expected convergence rate suggested by the corresponding Hilbert expansion.
More precisely, as shown in Appendix A, the kinetic equation
$\varepsilon\partial_{t}f^{\varepsilon}+\xi\cdot\nabla f^{\varepsilon}-\nabla(-\Delta)^{-\alpha}\rho^{\varepsilon}\cdot\nabla_{\xi}f^{\varepsilon}+\frac{1}{\varepsilon}\xi^{\perp}\cdot\nabla_{\xi}f^{\varepsilon}=\nabla_{\xi}\cdot(\nabla_{\xi}f^{\varepsilon}+\xi f^{\varepsilon})$
in two dimensions admits a formal expansion of the form
$f^{\varepsilon}=f_{0}+\varepsilon f_{1}+O(\varepsilon^{2}),$
where the leading-order term is given by
$f_{0}=M_{\rho,0},$
and the first-order corrector $f_{1}$ is given by
$f_{1}=-\left((\nabla\rho+\rho\nabla(-\Delta)^{-\alpha}\rho)^{\perp}\cdot\xi\right)M_{1,0},$
which encodes the effect of transport and nonlocal interaction after averaging along the fast rotational flow in velocity space.
At the macroscopic level, this first-order correction generates the effective flux
$m^{1}=-\nabla^{\perp}\rho-\rho\nabla^{\perp}(-\Delta)^{-\alpha}\rho,$
and the corresponding limit equation is
$\partial_{t}\rho-\nabla\cdot\bigl(\rho\nabla^{\perp}(-\Delta)^{-\alpha}\rho\bigr)=0.$
Equivalently, using $\nabla\cdot\nabla^{\perp}\phi=0$, one may write
$\partial_{t}\rho+\nabla^{\perp}(-\Delta)^{-\alpha}\rho\cdot\nabla\rho=0,$
which is the generalized surface quasi-geostrophic equation in active scalar form.
Therefore, at the level of the formal expansion, one expects
$f^{\varepsilon}-M_{\rho,0}=O(\varepsilon).$
Moreover, since the first-order corrector $f_{1}$ is odd in $\xi$, we have
$\int_{\mathbb{R}^{2}}f_{1}\,d\xi=0.$
Hence the first-order correction does not contribute to the density variable, and formally
$\rho^{\varepsilon}-\rho=O(\varepsilon^{2}).$
On the other hand, the momentum satisfies
$m^{\varepsilon}:=\int_{\mathbb{R}^{2}}\xi f^{\varepsilon}\,d\xi=\varepsilon\int_{\mathbb{R}^{2}}\xi f_{1}\,d\xi+O(\varepsilon^{2}),$
so that
$\frac{1}{\varepsilon}m^{\varepsilon}=-\nabla^{\perp}\rho-\rho\nabla^{\perp}(-\Delta)^{-\alpha}\rho+O(\varepsilon).$
Thus, as in the diffusive regime, the density variable gains one additional order compared to the kinetic distribution itself. This improvement is again due to the oddness of the first-order corrector in the velocity variable. We emphasize, however, that this is only a formal prediction of the asymptotic expansion; our rigorous estimates do not address this sharper $O(\varepsilon^{2})$ density rate.
Nevertheless, the first-order corrector can still be absorbed into a suitably shifted Maxwellian, which yields a more accurate kinetic approximation. Indeed, since
$f_{1}=-\left((\nabla\rho+\rho\nabla(-\Delta)^{-\alpha}\rho)^{\perp}\cdot\xi\right)M_{1,0}=-\xi\cdot\left(\nabla^{\perp}\log\rho+\nabla^{\perp}(-\Delta)^{-\alpha}\rho\right)M_{\rho,0},$
it is natural to introduce
${\rm u}_{\varepsilon}:=-\varepsilon\left(\nabla^{\perp}\log\rho+\nabla^{\perp}(-\Delta)^{-\alpha}\rho\right),$
so that
$f^{\varepsilon}=M_{\rho,0}+\xi\cdot{\rm u}_{\varepsilon}\,M_{\rho,0}+O(\varepsilon^{2}).$
Using the same first-order Taylor expansion of $M_{\rho,{\rm u}_{\varepsilon}}$ around ${\rm u}=0$ as in the diffusive case, we formally obtain
$M_{\rho,{\rm u}_{\varepsilon}}=M_{\rho,0}+\xi\cdot{\rm u}_{\varepsilon}\,M_{\rho,0}+O\bigl(\varepsilon^{2}(1+|\xi|^{2})M_{\rho,0}\bigr),$
and hence
$f^{\varepsilon}-M_{\rho,{\rm u}_{\varepsilon}}=O\bigl(\varepsilon^{2}(1+|\xi|^{2})M_{\rho,0}\bigr).$
Thus, the formal expansion suggests that the density variable satisfies
$\rho^{\varepsilon}-\rho=O(\varepsilon^{2}),$
while, at the kinetic level, the shifted local Maxwellian $M_{\rho,{\rm u}_{\varepsilon}}$ provides a sharper second-order approximation of $f^{\varepsilon}$.
Appendix C Propagation of regularity for the limit equations